OpenStudy (anonymous):

Solve for x: 2/3 (x - 2) = 4x

OpenStudy (kamibug):

\(\Huge\bf\color{seagreen}{Welcome~to~Openstudy!~:)}\)

OpenStudy (anonymous):

Welcome to open study!! :) X = - 2/5

OpenStudy (anonymous):

Thanks so much

OpenStudy (kamibug):

Distribute. 2/3x - 4/3 = 4x Combine Like Terms. -4/3 = 10/3x Divide both sides by 10/3. -12/30 = x Simplify. x = -2/5

OpenStudy (anonymous):

welcome, medal? :)

OpenStudy (mathstudent55):

\(\dfrac{2}{3}(x - 2) = 4x\) First, multiply both sides by 3/2 to get rid of the fraction: \(\color{red}{\dfrac{3}{2} \times}\dfrac{2}{3}(x - 2) = \color{red}{\dfrac{3}{2} \times}4x\) \(x - 2 = 6x\) Now subtract x from both sides: \(-2 = 5x\) Finally, divide both sides by 5: \(-\dfrac{2}{5} = x\)